OpenStudy (anonymous):

MATH QUESTION HELP! If the slope of the line that passes through the points (a, 0) and (1, -2) is 1/2, what is the value of a?

OpenStudy (anonymous):

The gradient is 1/2?

OpenStudy (anonymous):

The slope is 1/2. I'm not sure what you mean by gradient though.

OpenStudy (anonymous):

Gradient=slope

OpenStudy (anonymous):

The equation of the line is y=mx+c m=0.5 so y=0.5x+c To find c, insert (1,-2) into the equation (x=1, y=-2), as the line passes through this point. -2=0.5(1)+c c=-1.5 So y=0.5x-1.5 (a,0) insert y=0 0=0.5x-1.5 1.5=0.5x x=3

OpenStudy (anonymous):

*That method to find c works as c should satisfy the equation when any set of points (here (-2,1)) are put in. If x=3 when y=0, then a=3

OpenStudy (anonymous):

Do you want me to expand on anything you don't get?

OpenStudy (anonymous):

No I get it thanks a bunch! :)